Katie Bobinski, Abby Mitchell, Julie Van Saun – No Title

Your presentation to the class was very good. I understand that you might have been experiencing some time pressure since the other group that day went a little long. The main thing I would have liked to see there that you didn’t do was to go a bit deeper into the technical details of how you analyzed the survey data and how your conclusions there related to the other sections of the presentation.

The paper also has a number of good features on the content side. The reviews of past studies show some interesting points about whether environmental awareness can be fostered by educational efforts. You did some good work in analyzing the results from the other group’s survey. However, there are several major flaws too. The first is that the paper itself seems to have been “slapped together” quickly with not enough oversight from everyone involved. I know collaborating on a project like this might be a new experience for you and that you tried to address it by a division of labor. But coordinating how the sections fit together and how well the work stands as a single paper is important too.

For instance, the file I received had no title, and your names never appeared anywhere in it(!)

More importantly, there were a large number of citations in the text of the paper that did not correspond to any of the entries in the bibliography (works cited). Moreover, the citations don’t give page numbers or some other indication of where in the source you are taking the information from. I can guess how all that probably happened, but that just is not acceptable for a research paper like this. (Be aware, too, that in a lot of courses here at HC, such problems with citations would carry extremely steep penalties on the overall grade for the paper.)

Finally, there are a number of small but glaring language errors that recur throughout the paper. For instance, please figure out the difference between the words “since” and “sense” and the difference between “there” and “their” and learn which is appropriate where! These kinds of errors can destroy your credibility, even when you have something valuable to say.

Specific Comments

1) This first sentence does not really “work.” It is not the studies that are varying from person to person, it is the attitudes that vary.

2) Strictly speaking, you are claiming much too much about what you did in the second paragraph. How did your work “test whether or not a person’s scientific knowledge would affect their ability to comprehend that the environment is progressing toward an uninhabitable place?” It’s not the ability to comprehend that is the issue here. (The methods used in the Kahan, et. al. article and the HC survey don’t address that either.) The issue is whether people’s attitudes about environmental questions are correlated with their level of general scientific and quantitative knowledge.

2.5) Note that the Ballantyne, et. al. study is really addressing a somewhat different question than your main question. You could summarize their main question as “do environmental education programs increase communication about environmental issues between parents and children?” This might address how attitudes are formed, so it’s definitely related to your main issue, but you never really identify the way in which it is related.

3) Lots of sources cited here that don’t appear in the bibliography at the end!

4) This list of steps could apply to just about any scientific study. So I am not sure why you chose to include it here. Also, what does point (10) “defining its slope” mean?

5) Saying “the average level of agreement in this group was 25.87” does not convey much information to the reader unless you have described in more detail how the questions where constructed and how this level was computed.

6) The description of the Sebasto-Cavern study is not too good. First, explain what the acronyms NJSOC and NJMS mean. Then, I think you meant to say that since the study was not able to use a random sample, they used what is called a sample of convenience (you wrote “sample of convince” which is meaningless). Finally What does it mean to say the group of students was “expectable?” I have no idea what you were trying to say there.

7) “No true conclusion was found” is rather vague in describing the overall results of the Sebasto-Cavern study. It would be more precise to say that their data showed that the only case where there was a significant difference between the pre-test and post-test scores was one where the students had received extensive preparation and post-trip processing of the experience.

8) Your characterization of the Kahan, et. al. results – “people with higher scientific intellect had a lower perception of the risk of climate change” is not very precise. They did not try to measure “scientific intellect” (even assuming one could do that). They were concerned with scientific and mathematical literacy – the level of knowledge of basic facts and concepts.

9) Be more careful about how you interpret the p value of a statistical test of hypotheses. Saying a p value of .05 shows that “there was a 95% chance that math and science abilities did in fact affect a student’s views on environmental issues” is INCORRECT and a serious misunderstanding of the statistical method. The real meaning of p = .05 is that if the null hypothesis is true, then there is a 95% chance that we would see a “less extreme” value of the test statistic than the one we observed. The p value is essentially a measure of the strength of the evidence for rejecting or accepting the null hypothesis.

Final Project Grades

Presentation: A- (92)

Paper Content: B+ (88)

Paper Form: B (85)